OpenStudy (anonymous):

An offshore oil well is leaking oil and creating a circular oil slick. If the radius of the slick is growing at a rate of 5 miles/hour, find the rate at which the area is increasing when the radius is 7 miles. (The area of a disc of radius r is A = πr2.)

OpenStudy (anonymous):

\[\frac{ dx }{ dh }=5,\frac{ dA }{ dh }=?\]

OpenStudy (anonymous):

yes i just dont know how to set it up

OpenStudy (anonymous):

\[\frac{ dA }{ dt }=\frac{ dx }{ dt }*\frac{ dA }{ dx }\] \[\frac{ dA }{ dx }=2\pi r=2\pi(7)\]

OpenStudy (anonymous):

@Jonask is that the final answer?

OpenStudy (anonymous):

no you have to substitute in the chain rule formular

OpenStudy (anonymous):